OpenStudy (hawkfalcon):

A function f is defined on the interval [0,4], and its derivative is f'(x)=e^sinx - 2cos3x On what interval is f increasing? At what values of x does f have a local maxima? How many points of inflection does the graph of f have?

OpenStudy (hawkfalcon):

\[f'(x)=e ^{sinx} - 2\cos3x\]

OpenStudy (anonymous):

when is \[f'(x)>0\]solve

OpenStudy (anonymous):

\[e^{\sin x}>2\cos 3x\] \[\sin x>\ln(2\cos 3x)\]

OpenStudy (hawkfalcon):

Okay. O.o

OpenStudy (anonymous):